Askew Enterprises produces a product with fixed costs of $200,000 and variable cost of $9 per unit. The company desires to earn a $100,000 profit and believes it can sell 20,000 units of the product.

Required

  1. Based on this information, determine the target sales price.

  2. Assume a competitor is currently selling a similar product for $20 per unit. Explain how Askew can use target costing to maintain its desired profitability
Answer #1
a) Calculation of selling price at a target profit 100000
Fixed cost = 200000
Desired profit =100000
profit = quantity sold*selling price-variable cost-fixed cost
100000= 20000*selling price-(20000*9)-200000
selling price = 480000/20000 = 24
target selling price = 24
b) sale level
desired sale in unit = Fixed cost+required profit/contribution p.u
desired sale in unit = 200000+100000/(20-9)
desired sale in unit = 300000/11
desired sale in unit = 27273 units